Object A has an initial velocity of (10m/s, 0o). Object A has a mass of 5kg. Object B has a mass of 10kg and is initially at rest. After the collision, Object A heads off at an angle of 40o and Object B heads off at an angle of -20o. Calculate the velocity of each object after the collision.
1
Expert's answer
2020-04-15T10:24:46-0400

In case of collision initial and final momentum is same

initial momentum =mv=5"\times"10"i" + 0 = 50"i"

final momentum =(vacos40 + vbcos20)"i" + (vasin40 - vbsin20)"j"

for initial to be equal to final the "j" component must be zero

(vasin40 - vbsin20)=0

va(0.64)=vb(0.34)

vb=va(1.88) ..................(1)

now balancing the x component

vacos40 + vbcos20=50

va(0.76) + vb(0.94) =50

substituting the value of vb

va(0.76) + va(1.88)(0.94) = 50

2.5 va =50

va =20m/s

vb =10.64 m/s